Sarah is 10 years older than her sister Emma. At one time Sarah was three times as old as Emma. How old was Emma when Sarah was three times as old?
Answer by Alan3354(69443)

Sarah is 10 years older than her sister Emma. At one time Sarah was three times as old as Emma. How old was Emma when Sarah was three times as old?
----------
S = E + 10
Y = years before or after
S+Y = 3E+Y
E+10 + Y = 3E + Y
E = 5